OpenStudy (anonymous):

medal question Evaluate x/y for x = 3/10 and y = 4/5 A. 7/50 B. 7/15 C. 3/8 D. 12/50

OpenStudy (anonymous):

\[\frac{a/b}{c/d} = a/b \times d/c\]

OpenStudy (anonymous):

C.) 3/10 = .3 4/5 = .8 3/8

OpenStudy (anonymous):

simply the expression (2/9)^2

OpenStudy (anonymous):

I got 4/81

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Also, Baseballa's approach to the first question works very nicely in this case, although it doesn't really work in many other cases.

OpenStudy (anonymous):

what is the simplest form of the expression 8 [60 ÷ (4^2 - 2^2) - 2]

OpenStudy (anonymous):

@Indivicivet @Baseballa101

OpenStudy (anonymous):

I could solve it and I got 24

OpenStudy (anonymous):

which is a solution for the equation y = 4x + 3 A. (5, 17) B. (17, 5) C. (4, 19) D. (4, 16)